# FIT-GNN: Faster Inference Time for GNNs Using Coarsening

### v2

@@ Abstract
Scalability of Graph Neural Networks (GNNs) remains a significant challenge, particularly when dealing with large-scale graphs. To tackle this, coarsening-based methods are used to reduce the graph into a smaller graph, resulting in faster computation. Nonetheless, prior research has not adequately addressed the computational costs during the inference phase. This paper presents a novel approach to improve the scalability of GNNs by reducing computational burden during both training and inference phases. We demonstrate two different methods (Extra-Nodes and Cluster-Nodes). Our study also proposes a unique application of the coarsening algorithm for graph-level tasks, including graph classification and graph regression, which have not yet been explored. We conduct extensive experiments on multiple benchmark datasets in the order of $100K$ nodes to evaluate the performance of our approach. The results demonstrate that our method achieves competitive performance in tasks involving classification and regression on nodes and graphs, compared to traditional GNNs, while having single-node inference times that are orders of magnitude faster. Furthermore, our approach significantly reduces memory consumption, allowing training and inference on low-resource devices where traditional methods struggle.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work
[…]
images/img1
Figure 1: The figure shows how the coarsening algorithm reduces a graph $G$ with $15$ nodes to a coarsened graph $G^{\prime}$ with $3$ nodes. It also shows which nodes belong to which partition.
@@ 1 Introduction
Graph Neural Networks (GNNs) have proven to be a tool of surprising versatility and modeling capacity. However, there are still a few issues that limit their general applicability, the chief among these being the scalability of the GNN models. This scalability issue is problematic for both the training and the inference phases. Previous research has attempted to address this by reducing the graph into a smaller graph in order to reduce the computational cost. One such broad technique is graph coarsening methods, which judiciously remove nodes and edges from the graph. However, there are still certain points of weakness of these algorithms. For one, coarsening methods can potentially remove important node information. Secondly, and more pertinent to the current discussion, coarsening methods only reduce the training computation; they do not address the inference time. The next obvious question would be to ask why we cannot use coarsening methods for inference as well. Given a graph $G$, coarsening algorithms create a coarsened graph $G^{\prime}$. While $G^{\prime}$ is smaller than $G$, the nodes in $G^{\prime}$ are combinations of nodes in $G$, which can be both train and test nodes. As a result, the prediction per test node is not available. In another scenario, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes. Also, let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively. The coarsening algorithm will take the majority label and assign class 00 to $v^{\prime}$. Hence, the model will be trained to predict only the class 00, leading to discarding the quantification of the model’s performance on predicting less represented nodes. This is the research gap that we seek to address in this work. We use coarsening algorithms to partition the graph into a set of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. We show methods to train on these subgraphs that show equivalent results to non-coarsened baselines and sometimes perform better. This comes hand in hand with a significant reduction in inference time and memory consumption.
We demonstrate two different methods to append additional nodes in the subgraphs— Extra Nodes and Cluster Nodes that respectively append additional nodes to each subgraph, either by adding neighboring nodes, or by adding representational nodes corresponding a neighboring cluster. We theoretically characterize the upper bound of the coarsening ratio for which the time complexity of our approach is better than the naive uncoarsened approach. In node classification and node regression tasks, the baselines typically require the entire dataset during inference to get a prediction for a single node. In contrast, our method only needs to pass the subgraph to which the node belongs. Our results show up to $\mathbf{100\times}$ reduction in inference time and up to $\mathbf{70\times}$ less memory than the baseline while achieving equivalent or better performance.
@@ 2.1 Graph Notations
Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix. The feature of each node $v_{i}$ in the graph is the $i$th row vector in $X$ represented as $x_{i}\in\mathbb{R}^{d}$. Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges. Let us assume $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$. $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$. We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$.
@@ 2.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming node representations of its neighbors. Kipf and Welling (2017) introduced Graph Convolutional Network (GCN) as follows:
$X_{l+1}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X_{l}\mathcal{W}_{l})$ (1)
where $X_{l}$ is the node representation after $l$ layers of GCN, $\tilde{A}=A+I$, $\tilde{D}=D+I$, $I$ is an identity matrix of same dimension as $A$. $\mathcal{W}_{l}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
@@ 2.3 Graph Coarsening and Graph Partitioning
Loukas (2018) discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $k=n\times r$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$.
For the coarsened graph, the adjacency matrix $A^{\prime}=P^{T}AP$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. Huang et al. (2021) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{j,j}=|C_{j}|,j=1,2,\dots,k$.
Along with a coarsened graph, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$. The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$. Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$. $A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
3 Related Work
The goal of reducing computation costs has been tackled via different methodologies. Research like Chiang et al. (2019) and Zeng et al. (2020) use a subgraph sampling-based approach. They sample a small subgraph at each training iteration and train the GNN. Other techniques like layer-wise sampling along with mini-batch training have been studied by  Chen et al. (2018); Chen and Zhu (2018); Cong et al. (2020); Zou et al. (2019). Fahrbach et al. (2020) introduced a new coarsening algorithm based on Schur complements to generate embeddings for vertices in a large graph. Huang et al. (2021) introduced using GNN for a coarsened graph. The motivation was to utilize graph coarsening algorithms to reduce the number of nodes and edges in the graph, resulting in less training time. This approach helps to train GNN for graphs with a large number of nodes. However, inference time is not reduced because the nodes cannot be removed during inference. This training process also caused the removal of important node-level information. Later, Xue et al. (2022) explored shifting the training process from graph-level training to subgraph-level. Coarsening algorithms were used to split the graph into $k-$subgraphs, resulting in the removal of a few edges. Further, to speed up the training process, they propose a multi-device training approach where subgraphs are sent to different devices. This results in a faster training process. Kumar et al. (2023) further proposed the usage of feature-based coarsening algorithms.
4 Methodology
In Huang et al. (2021), a graph $G=(V,E,X,W)$ is reduced to $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ using a partition matrix $P$ generated from the coarsening algorithm. Here $X^{\prime}=P^{T}X$. The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task. While using $\arg\max$, some essential information is lost. Therefore, we follow a subgraph-level training process where we don’t remove these label information. To explore another node-level task, such as node regression, functions such as the mean of regression targets wouldn’t be appropriate because they lose the variance of the targets. Hence, our goal is to use subgraphs for training and inference purposes.
To tackle the issue of loss of information corresponding to the nodes (and their edges) present at the periphery of the subgraphs after partitioning the graph, we append nodes in two ways:
@@ 4 Methodology
Figure 2: This figure shows the methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
$\bullet$ Extra Nodes: Adding the $1$-hop neighbouring nodes in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$.
$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}$
However, we don’t perform backpropagation of the prediction on $\mathcal{E}_{G_{i}}$. Xue et al. (2022) introduced this approach to reduce information loss after partitioning. An edge between two Extra Nodes in $\mathcal{E}_{G_{i}}$ is added if they were connected in $G$.
Algorithm 1 Node-Model$(L,A,D,X)$
1:Number of Layers $L$; $A$; $D$; $X_{0}=X$;
2:for $i=1$ to $L$ do
[…]
4:end for
5:$Z=X_{L}\mathcal{W}_{L}$
6:return $Z$
$\bullet$ Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for GNN models with a large number of layers, information loss will still persist. Instead of adding the neighboring nodes, we will add a node, namely Cluster Nodes, representing the neighboring cluster, denoted by $\mathcal{C}_{G_{i}}$.
$\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$
Liu et al. (2024) introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method after adding these Cluster Nodes and cross-cluster edges. In our work, we add cross-cluster edges; however, we do not sample from the constructed set of subgraphs.
Figure 2 illustrates appending these additional nodes using these approaches. The dashed line shows which nodes are added as part of these additional nodes. Predictions on these newly appended nodes will not contribute to loss calculation and weight update.
Lemma 1.
@@
There are $2$ sets of nodes in $G_{i}$ as follows:
  • The set $S_{1}$ of nodes with 1-hop neighbours in $G_{i}$.
  • The set $S_{2}$ of nodes with not all 1-hop neighbors in $G_{i}$
Now, |$\begin{split}\textup{Now, }\left|\bigcup_{v\in S_{1}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\right|=0\\ \textup{Also, }\mathcal{E}_{G_{i}}=\bigcup_{v\in S_{2}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\\ \end{split}$
$\Rightarrow\left|\mathcal{E}_{G_{i}}\right|=\mathcal{I}^{1}_{i}$
Hence when Extra Nodes is used, 1 Layer GNN model cannot distinguish $G$ and $\mathcal{G}_{s}$. ∎
@@
1:$G=(V,E,X)$; Labels $Y$; Model $M$; Loss $\ell$;
@@
4:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$. Masks denote the nodes that are used for training;
5:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
6:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
7:$O\leftarrow[]$; $Y\leftarrow[]$;
8:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
9:     $O_{i}=M(A_{i},D_{i},X_{i})$;
10:     $O.append(O_{i}[mask_{i}])$;
11:     $Y.append(Y_{i}[mask_{i}])$;
12:end for
13:$Loss=\ell(O,Y)$;
14:Train $M$ to minimize $Loss$;
Using Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
$\bullet$ The number of nodes added to a subgraph via Cluster Nodes is less than or equal to the number of nodes added via the Extra Nodes approach. This is because, using Extra Nodes, we add all the neighboring nodes. However, we only add one representative node for each neighboring cluster using Cluster Nodes. As the GNN propagation depends on the number of nodes, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
$\bullet$ As mentioned before, the Extra Nodes approach reduces the information loss due to partitioning. However, with multiple layers of GNN, longer dependencies are not captured. In contrast, the Cluster Nodes approach overcomes this by computing a single Cluster Nodes’s representation as a combination of features of all the nodes present in the corresponding cluster. This results in sharing further node information in $1$-hop. Additionally, the transfer of information from one cluster to another is captured.

### v3

# FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening

@@ Abstract
Scalability of Graph Neural Networks (GNNs) remains a significant challenge. To tackle this, methods like coarsening, condensation, and computation trees are used to train on a smaller graph, resulting in faster computation. Nonetheless, prior research has not adequately addressed the computational costs during the inference phase. This paper presents a novel approach to improve the scalability of GNNs by reducing computational burden during the inference phase using graph coarsening. We demonstrate two different methods – Extra Nodes and Cluster Nodes. Our study extends the application of graph coarsening for graph-level tasks, including graph classification and graph regression. We conduct extensive experiments on multiple benchmark datasets to evaluate the performance of our approach. Our results show that the proposed method achieves orders of magnitude improvements in single-node inference time compared to traditional approaches. Furthermore, it significantly reduces memory consumption for node and graph classification and regression tasks, enabling efficient training and inference on low-resource devices where conventional methods are impractical. Notably, these computational advantages are achieved while maintaining competitive performance relative to baseline models.
@@ 1 Introduction
Graph Neural Networks (GNNs) have demonstrated remarkable versatility and modeling capabilities. However, several challenges still hinder their widespread applicability, with scalability being the most significant concern. This scalability issue affects both the training and inference phases. Previous research has sought to mitigate the higher training computation by using three broad categories of methods: 1. Training every iteration on the sampled subgraph 2. Training on a coarsened graph generated from a coarsening algorithm 3. Training on a synthetic graph that mimics the original graph . Unfortunately, these smaller graphs can only be utilized for training purposes and have not alleviated the computational burden during inference since the method mentioned above requires the whole graph during the inference phase. Consequently, for larger graphs, not only do inference times increase, but memory consumption may also exceed limitations.
To mitigate the high computational cost associated with the inference phase, we decompose the original graph $G$ into a collection of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. This transformation yields substantial improvements in inference efficiency, achieving up to a $\mathbf{100\times}$ speedup for single-node prediction. Moreover, it significantly reduces memory consumption by as much as $\mathbf{100\times}$ —thereby enabling scalable inference on large-scale graphs comprising millions of edges. We provide a theoretical characterization of the upper bound on the coarsening ratio under which our method’s time and space complexity remain superior to that of the standard approach. Notably, we observe that conventional methods often fail to perform inference on large graphs due to memory constraints, whereas our approach remains effective and scalable in such scenarios.
While partitioning the graphs into subgraphs seems to be a naive transformation, it comes with the cost of information loss corresponding to the edges removed. We overcome this loss by appending additional nodes in the subgraphs using two different methods – Extra Nodes and Cluster Nodes that respectively append neighboring nodes or representational nodes corresponding to neighboring clusters. These additional nodes resulted in equivalent or better performance in terms of accuracy and error compared to baselines. Our contributions are summarized as follows:
  1. 1.
    We reduce inference time and memory by applying graph coarsening to partition graphs into subgraphs, mitigating information loss through Extra Node and Cluster Node strategies.
  2. 2.
    We provide a theoretical analysis of our approach, establishing a lower inference computational complexity compared to baselines.
  3. 3.
    Our method scales well to extremely large graphs that baselines fail to process, while consistently delivering strong performance on 13 real-world datasets.
Figure 1: The figure shows the overall pipeline of our proposed method and its comparison with traditional training and inference of GNNs. This pipeline is made for node-level tasks.
2 Related Work
The aim of reducing computational costs has been approached through various methods. Research by Chiang et al. [1] and Zeng et al. [2] uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training. Other methods, including layer-wise sampling and mini-batch training, have been explored by Chen and Zhu [3], Chen et al. [4], Cong et al. [5], Zou et al. [6]. Additionally, Fahrbach et al. [7] introduced a coarsening algorithm based on Schur complements to create embeddings for vertices in large graphs, while Huang et al. [8] (SGGC) applied GNNs to coarsened graphs, reducing nodes and edges to lower training time. Kumar et al. [9] proposed feature-based coarsening algorithms. Jin et al. [10] (GCOND), Wang et al. [11] (GC-SNTK) presented algorithms to generate a synthetic graph that mimics the original graph. However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden. Gupta et al. [12] (BONSAI) addressed this issue by looking into computational trees and training on just the relevant computational trees. However, these approaches do not decrease inference time since the inference phase takes the whole graph as input. Later, Xue et al. [13] shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach. In the graph-level task, Jin et al. [14] (DOSCOND) and Xu et al. [15] (KIDD) tackled the graph classification problem following up on the approach of GCOND and GC-SNTK. However, the graph regression problem was not explored in this direction.
@@ 3.1 Graph Notations
Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix. Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges. Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$. $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$. We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$.
@@ 3.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors. Kipf and Welling [16] introduced Graph Convolutional Network (GCN) as follows:
$X^{(l+1)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(l)}\mathcal{W}^{(l)})$ (1)
where $X^{(l)}$ is the node representation after $l$ layers, $\tilde{A}=A+I$, $\tilde{D}=D+I$, $I$ is the identity matrix of same dimension as $A$. $\mathcal{W}^{(l)}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
@@ 3.3 Graph Coarsening and Graph Partitioning
Loukas [17] discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $k=n\times r$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$. $A^{\prime}=P^{T}AP$ is the adjacency matrix of $G^{\prime}$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. SGGC used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
Along with $G^{\prime}$, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$. The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$. Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$. $A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
@@
Figure 2: Methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
4 Methodology
In SGGC, a graph $G$ is reduced to $G^{\prime}$ using a partition matrix $P$ generated from the coarsening algorithm. Here $X^{\prime}=P^{T}X$. The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task. Training is carried out on the coarsened graph $G^{\prime}$ using $Y^{\prime}$ as the label information. While this approach is novel, essential label information is lost. For instance, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes. Let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively. $\arg\max$ will take the majority label and assign class 0 to $v^{\prime}$. Hence, the model will be trained to predict only the class 0, leading to the discarding of the quantification of the model’s performance on predicting less represented nodes.
Similarly, to explore another node-level task, such as node regression, aggregation functions such as the mean of regression targets wouldn’t be appropriate because they lose the variance of the targets. To mitigate label information loss, we use original label information without any aggregation function like $\arg\max$ or mean, thus motivating us to follow a subgraph-level training process.
However, partitioning the graph into subgraphs leads to a loss of neighborhood information at the periphery of subgraphs. To tackle this, we append nodes in two ways after partitioning:
  • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$. Xue et al. [13] introduced this approach to reduce information loss after partitioning. We add an edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$.
  • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist. Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$. Liu et al. [18] introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph. In our work, we add cross-cluster edges but do not sample from the constructed set of subgraphs.
$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\quad\text{and}\quad\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$ (2)
Figure 2 illustrates appending these additional nodes using these approaches. The dashed line shows which nodes are added as part of these additional nodes. For subgraph-level training, specifically for node-level tasks, the newly appended nodes do not contribute to the weight update as the loss is not backpropagated based on the predictions of these nodes. This is done since the appended nodes might contain information from the test nodes. Therefore, a boolean array $mask_{i}$ is created for each subgraph $G_{i}$ such that $True$ is set for nodes that actually belong to the subgraph $G_{i}$ (not as an Extra Node or Cluster Node) and are training nodes; otherwise, $False$.
Lemma 4.1.
@@ 4 Methodology
In Appendix A.1. ∎
Using the Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
  • From Equation 2, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$ This is because, using Extra Nodes, we add all the neighboring nodes. However, using Cluster Nodes, we only add one representative node for each neighboring cluster. Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
  • Lemma 4.1 shows that the Extra Nodes approach reduces the information loss due to partitioning for $1$ Layer of GNN. However, with multiple layers of GNN, longer dependencies are not captured. In contrast, the Cluster Nodes approach overcomes this by computing a single Cluster Node’s representation as a combination of features of all the nodes present in the corresponding cluster. This results in sharing further node information in $1$-hop. Additionally, the transfer of information from one cluster to another is captured.
@@
1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
@@
4:if Append additional nodes then
[…]
6:end if
7:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
8:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
9:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
[…]
11:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
12:  $O_{i}=M(L,A_{i},D_{i},X_{i})$; $\quad\#$ Algorithm 4
13:  $O.append(O_{i}[mask_{i}])$;
14:  $Y.append(Y_{i}[mask_{i}])$;
15:end for
[…]
17:Train $M$ to minimize LossLoss;